### 17 Dec 2023 v2

# SUREL+: Moving from Walks to Sets for Scalable Subgraph-based Graph Representation Learning

Abstract.
Subgraph-based graph representation learning (SGRL) has recently emerged as a powerful tool in many prediction tasks on graphs due to its advantages in model expressiveness and generalization ability. Most previous SGRL models face computational issues associated with the high cost of subgraph extraction for each training or test query. Recently, SUREL was proposed to accelerate SGRL, which samples random walks offline and joins these walks online as a proxy of subgraphs for representation learning. Thanks to the reusability of sampled walks across different queries, SUREL achieves state-of-the-art performance in terms of scalability and prediction accuracy. However, SUREL still suffers from high computational overhead caused by node redundancy in sampled walks. In this work, we propose a novel framework SUREL+ that upgrades SUREL by using node sets instead of walks to represent subgraphs. This set-based representation avoids repeated nodes by definition, but node sets can be irregular in size. To address this issue, we design a customized sparse data structure to efficiently store and index node sets, and provide a specialized operator to join them in parallel batches. SUREL+ is modularized to support multiple types of set samplers, structural features, and neural encoders to complement the structure information loss after the reduction from walks to sets. Extensive experiments have been performed to validate SUREL+ in the prediction tasks of links, relation types, and higher-order patterns. SUREL+ achieves 3-11$\times$ speedups of SUREL while maintaining comparable or even better prediction performance; compared to other SGRL baselines, SUREL+ achieves $\sim$20$\times$ speedups and significantly improves the prediction accuracy.
PVLDB Reference Format:
PVLDB, 16(11): 2939-2948, 2023.
doi:10.14778/3611479.3611499 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 16, No. 11 ISSN 2150-8097.
doi:10.14778/3611479.3611499
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/Graph-COM/SUREL_Plus.
1. Introduction
Graphs are widely used to model interactions in natural sciences and relationships in social life (Koller et al., 2007; Jumper et al., 2021). Graph-structured data in the real world are highly irregular and often large-scale. To solve inference tasks on graphs, graph representation learning (GRL) that studies quantitative representations of graph-structured data has attracted much attention (Hamilton et al., 2017b; Hamilton, 2020; Wu et al., 2022). Recently, subgraph-based GRL (SGRL) has become an important research direction for researchers studying GRL algorithms and systems, as it achieves far better prediction performance than other approaches in many GRL tasks, especially those involving a set of nodes. Given a set of nodes of interest, namely a queried node set, SGRL models such as SEAL (Zhang and Chen, 2018; Zhang et al., 2021), GraIL (Teru et al., 2020), and SubGNN (Alsentzer et al., 2020) first extract a subgraph around the queried node set (called query-induced subgraph), and then use neural networks to encode extracted subgraphs for prediction. Extensive work shows that SGRL models are more robust (Zeng et al., 2021) and more expressive (Bouritsas et al., 2022; Frasca et al., 2022); while canonical graph neural networks (GNNs) including GCN (Kipf and Welling, 2017) and GraphSAGE (Hamilton et al., 2017a) usually fail to make accurate predictions, due to their limited expressive power (Zhang et al., 2021; Garg et al., 2020; Chen et al., 2020), incapability of capturing intra-node distance information (Srinivasan and Ribeiro, 2020; Li et al., 2020), and improper entanglement between receptive field size and model depth (Huang and Zitnik, 2020; Zeng et al., 2021; Yin et al., 2022). An example in Fig. 2 illustrates how SGRL works for link prediction and demonstrates its advantages over GNNs. Here, canonical GNNs generate and aggregate node-wise representations to predict links, which would map structurally symmetric nodes into the same representation and lead to the ambiguity issue (Xu et al., 2019; Zhang et al., 2021). So far, the advantages of SGRL methods have been proved in many applications, such as link and relation prediction (Zhang and Chen, 2018; Zhang et al., 2021; Teru et al., 2020), higher-order pattern prediction (Meng et al., 2018; Liu et al., 2022), temporal network modeling (Wang et al., 2021), recommender systems (Zhang and Chen, 2020), anomaly detection (Alsentzer et al., 2020; Cai et al., 2021), graph meta-learning (Huang and Zitnik, 2020), subgraph matching (Liu et al., 2020; Lou et al., 2020), and molecular/protein research in life sciences (Wang and Zhang, 2021).
Figure 1. GNNs cannot correctly predict whether $x$ is more likely linked with $y$ or $z$: without distinct features, $y$ and $z$ have the same node representations. The representations based on one-hop neighbors (query-induced subgraph) are more expressive in distinguishing these two node pairs.
Figure 1. GNNs cannot correctly predict whether $x$ is more likely linked with $y$ or $z$: without distinct features, $y$ and $z$ have the same node representations. The representations based on one-hop neighbors (query-induced subgraph) are more expressive in distinguishing these two node pairs.
Figure 2. Overview of SUREL+: a side-by-side comparison with previous SGRL models. For a query, SGRL first prepares the input subgraph paired with structural features, which are then fed into a neural encoder to obtain embeddings for prediction. SUREL+ samples node sets as input, while SEAL (Zhang and Chen, 2018; Zhang et al., 2021) extracts induced subgraphs and SUREL (Yin et al., 2022) samples walks. To compensate for structure loss by reducing subgraphs to node sets, SUREL+ provides various types of set samplers, structure encoders, and set neural encoders. To better serve set-based representations, SUREL+ designs a customized sparse data structure SpG to store sampled node sets with fast access, and supports their online joining in parallel via a sparse join operator SpJoin.
Albeit with multiple benefits of its algorithm, SGRL methods currently face two major computational challenges: (1) Query Dependency. A subgraph must be extracted for each queried node set, which is not reusable across different queries, and cannot be preprocessed if the query is unknown; (2) Irregularity. The extracted subgraphs are irregularly sized, resulting in poor batch processing and load-balancing performance. As shown in Fig. 3 (a), subgraph extraction in SEAL (Zhang and Chen, 2018; Zhang et al., 2021) is prohibitively slow to be deployed in practice. This inspired recent work on dedicated hardware acceleration for subgraph extraction (DGL, 2022; PyG, 2022). However, how to fundamentally improve the scalability and efficiency of SGRL methods remains largely unexplored.
Figure 3. (a) Subgraph extraction in SEAL (Zhang and Chen, 2018; Zhang et al., 2021) has much higher complexity than other simplified form samplers. The size of induced subgraphs is irregular, growing exponentially with the number of extracted hops (Chen et al., 2018; Zeng et al., 2021). (b) Breaking subgraphs into reusable walks reduces complexity but faces the issue of heavy node duplicates.
SUREL (Yin et al., 2022) is the state-of-the-art (SOTA) framework that applies algorithm and system co-design to implement SGRL. It employs the join of node-level sampled walks to represent query-specific subgraphs. Specifically, SUREL treats each node as a seed and runs multiple random walks from the seed on the graph offline. Given a queried node set, SUREL online joins and encodes the sampled walks for all queried nodes for prediction. The join operation builds the connection between sampled walks of queried nodes so that their joint walks can function as the query-induced subgraph. To compensate for the structure loss by representing subgraphs in walks, a structural feature termed relative position encoding (RPE) is adopted to record the position of each distinct node in sampled walks relative to the seed. The RPE is pre-computed offline and attached to sampled walks before being fed into neural networks (NNs) to make predictions. In SUREL, sampled walks from one seed can be reused for multiple queries whenever that seed node is involved. SUREL significantly improves the scalability of SGRL by this walk-sharing mechanism and benefits from the regularity of walks, enabling highly parallel walk sampling and online joining through a dedicated system design. However, SUREL still faces several inherent drawbacks of the walk-based representation, namely high node redundancy in sampled walks (over 55% duplication, see Fig. 3 (b)). This further raises the following issues: (1) extra space for hosting walks and positional encoding in memory, (2) extra time for operations on repeated nodes in subsequent routines of walk joining and NN-based encoding, and (3) high workload of data transfer between CPU and GPU.
In this work, we upgrade SUREL and develop a novel framework SUREL+ that again benefits from algorithm-system co-design. The core concept of SUREL+ is simple: instead of using walks to represent subgraphs, we now employ node sets, thereby obviating node duplication. However, this new idea based on node sets introduces several algorithm and system design difficulties. From an algorithmic perspective, as explored in this study, the transition from induced subgraphs to walks and then to node sets results in a considerable loss of structural information. Therefore, the first priority is to develop a method that can compensate for structural losses while maintaining performance. On the system side, SUREL (Yin et al., 2022) utilizes walks, which can be easily stored and processed in an aligned format by controlling the length and number of sampled walks. In contrast, node sets are irregular in size, creating difficulties for efficient storage and fast access. To sum up, how to coordinate the designs of both sides constitutes the primary challenge of this work.
SUREL+ tackles the above challenges throughout its entire pipeline. To avoid node duplication in walk sampling, SUREL+ only keeps unique nodes from each seed node’s sampled neighborhood during preprocessing. To compensate for structure information loss, SUREL+ incorporates various types of set samplers and structure encoders to preserve local structures: set samplers employ different graph metrics to measure node importance and determine sampling rules; structure encoders support landing probabilities of random walks (Li et al., 2019), shortest path distances, and personalized PageRank scores (Jeh and Widom, 2003), covering most of the structural features used by previous SGRL models (Zhang and Chen, 2018; Zhang et al., 2021; Li et al., 2020; Teru et al., 2020; Yin et al., 2022). Furthermore, SUREL+ designs a customized sparse data structure, namely SpG, which can efficiently store sampled node sets and achieve fast access. During training and inference, SUREL+ implements a sparse operator SpJoin to perform join operations on the sampled node sets and their associated structural features of all queried nodes, acting as the query-induced subgraph for prediction. In addition, to capture diverse levels of interactions between node and structural features, SUREL+ supports multiple set neural encoders, such as multi-linear perception (MLP) with mean pooling, set attention (Veličković et al., 2018) and LSTM (Hamilton et al., 2017a) that ensure sufficient expressive power and consistent performance across various types of SGRL tasks.
Overall, our contributions can be summarized as follows:
  • Algorithm: SUREL+ is a novel SGRL framework (open source), that utilizes reusable node sets and associates them with various types of structural features to represent query-induced subgraphs via online joining. Adopting sets in SUREL+ greatly reduces memory and computation costs without degrading prediction performance compared with the SOTA baselines.
  • System: SUREL+ designs a customized sparse data structure SpG and a sparse join operator SpJoin to support efficient storage and processing of node sets, which achieves much better efficiency and scalability than previous SGRL methods.
  • We conduct extensive experiments on 9 real-world graphs, with millions/billions of nodes/edges, and demonstrate the advantages of SUREL+ in link/relation-type/motif prediction tasks. SUREL+ is 3-11$\times$ faster than the current SOTA SGRL method SUREL while maintaining comparable or even better prediction accuracy. SUREL+ also achieves $\sim$20$\times$ speedups with substantial accuracy improvements over other SGRL baselines.
2. Preliminaries
2.1. Notations and Relevant Definitions in SGRL
Let $\mathcal{G}(\mathcal{V},\mathcal{E},X)$ be an attributed graph with node set $\mathcal{V}=\{1,2,...,n\}$ and edge set $\mathcal{E}$, where $X\in\mathbb{R}^{n\times d}$ denotes node attributes with $d$-dimension. A query $Q\subset\mathcal{V}$ is a node set of interest for a certain type of task. We denote the subgraph induced by query $Q$ as $\mathcal{G}_{Q}$ and the node-induced subgraph as $\mathcal{G}_{u}$, where induced subgraphs are typically within a small number of hops.
Definition 2.0 (Subgraph-based Graph Representation Learning (SGRL)).
Given a query $Q$ of node set over graph $\mathcal{G}$, SGRL aims to learn a representation of the query-induced subgraph $\mathcal{G}_{Q}$ to make prediction $f(\mathcal{G}_{Q})$. ()($\cdot$)( ⋅ ) is usually a neural network. SGRL tasks come with some labeled queries $\{(Q_{i},y_{i})\}_{i=1}^{L}$ for supervision (positive samples) and other unlabeled queries $\{Q_{i}\}_{i=L+1}^{L+N}$ for inference.
Examples of SGRL Tasks Link prediction seeks to estimate the likelihood of a link between two nodes in a given graph, where a query $Q$ corresponds to a node pair. It can be further generalized to predict links with types over heterogeneous graphs (Teru et al., 2020) or to predict blood vessels (Paetzold et al., 2021) and chemical bonds (Jumper et al., 2021) in domain-specific graphs. Tasks beyond pairwise relations are named higher-order pattern prediction, where a query $Q$ consists of three or more nodes. In this work, we consider that given partially observed pairwise relations among queried nodes in $Q$, whether these queried nodes will establish certain full higher-order relation of interest (Srinivasan et al., 2021; Liu et al., 2022).
Review of SGRL Methods The current SGRL pipeline mainly has three parts, as shown in the Algorithm Design section of Fig. 2: subgraph preparation, structural feature construction, and neural encoder to obtain subgraph embeddings. Classical SGRL models often group query-dependent parts together (e.g., SEAL (Zhang and Chen, 2018; Zhang et al., 2021) couples subgraph extraction with labeling trick (Zhang et al., 2021)), and then apply neural encoder on extracted subgraphs with attached structural features for prediction. However, such coupling is expensive and makes the computed intermediate results not reusable across queries, which motivates recent SGRL methods to decouple them. SUREL (Yin et al., 2022) substitutes explicit subgraph extraction with the online joining of multiple walks pre-sampled offline associated with position encoding defined on sampled walks as structural features. Sampled walks and positional encoding are node-level and can be reused to serve multiple queries, which greatly improves efficiency. Lastly, neural networks are applied to encode and aggregate the embedding of joint walks for prediction.
2.2. Related Works
Scalable SGRL Design. Recent works on SGRL models have primarily focused on efficient subgraph extraction. Various techniques have been proposed, including PPR-based (Bojchevski et al., 2020; Zeng et al., 2021) and random walk-based (Yin et al., 2022) subgraph samplers, node neighborhood sampling through CUDA kernel (DGL,  (DGL, 2022)) and tensor operations (PyG, (PyG, 2022)). Some frameworks have customized data structures to better support subgraph operations and gain higher throughput, such as associative arrays in SUREL (Yin et al., 2022), temporal-CSR in TGL (Zhou et al., 2022), and GPU-orientated dictionary in NAT (Luo and Li, 2022). To achieve scalable modeling design, GDGNN (Kong et al., 2022) utilizes node representations along the geodesic path between queried nodes for prediction, partially decoupling structural feature construction from subgraph extraction. BUDDY (Chamberlain et al., 2023) employs subgraph sketches to avoid explicitly constructing subgraphs for link prediction. However, these works either focus on specific aspects of the SGRL scalability issue, namely extraction, storage, and feature construction bottlenecks, or are limited to one type of SGRL tasks. In contrast, SUREL+ provides a comprehensive co-design approach in scalable sampling, efficient storage, and expressive modeling, offering a general subgraph-based framework for scalable SGRL.
3. The framework of SUREL+
This section introduces SUREL+, whose key concept is to offline sample node sets around seed nodes in the graph, which can be joined online as a proxy of query-induced subgraph for representation learning. This approach keeps only distinct nodes in the sampled set for reuse in different queries, effectively addressing memory and computation concerns of node duplication in existing walk-based representations adopted by SUREL (Yin et al., 2022). SUREL+ features a modular design that supports various set samplers and structure encoders, and different set neural encoders can be paired with diverse structural features to compensate for the structure information loss after reducing subgraphs to node sets. Furthermore, SUREL+ introduces a customized sparse data structure SpG and an arithmetic operator SpJoin to store node sets and perform their online joins efficiently. Fig. 2 summarizes and compares SUREL+ and current SGRL models. The following subsections describe these modules in detail.
3.1. Set Samplers and Structure Encoders
SUREL+ uses set samplers to sample a set of nodes from the neighborhood of each seed node in the graph and calls structure encoders to construct structural features. Both operations are executed offline: the former is primarily for computational benefits, while the latter is to offset the structure loss of node sets reducing from subgraphs (adopted by SEAL (Zhang and Chen, 2018; Zhang et al., 2021)) or walks (adopted by SUREL (Yin et al., 2022)). Conceptually, SUREL+ represents the node-induced subgraph $\mathcal{G}_{u}$ via a combination of (1) a node set $\mathcal{S}_{u}$ comprising unique nodes sampled from the neighborhood of node $u$ and (2) the associated structural features $\mathcal{Z}_{u}$ reflects the position of sampled nodes in $\mathcal{G}_{u}$.
Set Samplers Two types of set samplers are adopted. The first type, named Walk-based Sampler, is to sample short-step random walks and eliminate repeated nodes during sampling. The second type, named Metric-based Sampler, is based on more principled graph metrics that measure the proximity between neighboring nodes and the seed node, such as personalized PageRank (PPR) scores (Jeh and Widom, 2003) or short path distances. Specifically, the walk-based sampler runs $M$-many $m$-step random walks, starting from each seed $u$ in parallel on the graph $\mathcal{G}$, and then puts only distinct nodes in sampled walks into set $\mathcal{S}_{u}$. The metric-based sampler, taking PPR-based (Bojchevski et al., 2020) as an example, first runs the push-flow algorithm (Andersen et al., 2006) to obtain an approximation of the PPR vector for each seed $u$, and then selects the top-$K$ nodes with the highest PPR scores into set $\mathcal{S}_{u}$. Mathematically, PPR scores are convergent landing probabilities of seeded random walks that reach infinite steps. Therefore, these two samplers complement each other by leveraging either more local or global structures of the graph. We use hyper-parameters $M$, $m$ to control random walks, and $K$ to control metric-based samplers, which are all set as some constants in practice. The complexity of the above offline sampling procedures is $O(|\mathcal{V}|)$.
Structure Encoders The structure encoder is to construct structural features $\mathcal{Z}_{u,x}\in\mathbb{R}^{k}$ for each node $x$ in the sampled node set $\mathcal{S}_{u}$. These features prove to be crucial for inference tasks on graphs involving multiple nodes (Zhang et al., 2021) and can be conceptually understood as anchoring sampled node $x$ in the seed $u$’s neighborhood. One possible choice is landing probabilities of random walk (Li et al., 2019, 2020; Yin et al., 2022): each element $\mathcal{Z}_{u,x}[i]$ stores the counts of node $x$ landed at step $i$ of all walks rooted at the seed node $u$ divided by the total number of walks performed by the sampler. Landing probabilities (LPs) can be computed along with walk sampling. Another option is the shortest path distance (SPD) between $x$ and $u$ (Zhang and Chen, 2018; Li et al., 2020; Zhang et al., 2021), which records their relative distance in terms of reachability. PPR scores (Jeh and Widom, 2003) is another popular structural feature and can be naturally obtained by running a PPR-based sampler. Later, we denote the collection of structural features for all nodes in $\mathcal{S}_{u}$ as $\mathcal{Z}_{u}=\{\mathcal{Z}_{u,x}|x\in\mathcal{S}_{u}\}$.
Figure 4. Node set $\mathcal{S}_{u}$ and its associated structural features $\mathcal{Z}_{u}$ stored in SpG. It contains three arrays to store the node set size, its elements, and associated structural features (with optional two-level indexing). Here, $D_{\text{SF}}$ shows the landing counts of nodes at different steps in sampled walks as an example, which can be normalized later to landing probabilities as structural features.
3.2. Set-based Storage - SpG
Node-set-based representations have advantages in terms of flexibility and compactness. However, the uneven sizes of sampled node sets pose great challenges to their storage and fast access. Note that these node sets are frequently visited in subsequent online phases for inference. To overcome these obstacles, SUREL+ designs a customized compressed sparse row (CSR) format called SpG, which reorganizes the storage of node sets and their structural features in a memory-efficient manner, as depicted in Fig. 4. Specifically, the node set $\mathcal{S}_{u}$ and its structural features $\mathcal{Z}_{u}$ are stored as a row of SpG, denoted as $\texttt{SpG}[u,:]$. Multiple node sets and their associated structural features are consolidated into three contiguous arrays:
  • indptr $\delta\in\mathbb{Z}^{n+1}$, an integer array tracks the starting index of each stored node set (row). It records the cumulative sum of the sizes of all node sets $\mathcal{S}_{u},~{}\forall u\in\mathcal{V}$, e.g., $\delta_{u+1}=\delta_{u}+|\mathcal{S}_{u}|$, where $|\mathcal{S}_{u}|$ represents the size of the set $\mathcal{S}_{u}$. The total number of sampled nodes stored in SpG is $\delta_{n+1}$;
  • indices $I\in\mathbb{Z}^{\delta_{n+1}}$, a coalesce array of all node sets $\mathcal{S}_{u},~{}\forall u\in\mathcal{V}$. The segment $I[\delta_{u}:\delta_{u+1}]$ corresponds to the indices of sampled nodes of $\mathcal{S}_{u}$ stored in sorted order. This ordering is particularly useful for speeding up the join operation discussed in Sec. 3.3.
  • SFptr $D\in\mathbb{R}^{\delta_{n+1}}$, a pointer array contains the indices of the structural features $\mathcal{Z}_{u}$ stored in the array $D_{\text{SF}}$. The purpose of $D_{\text{SF}}$ is to eliminate duplicate structural features that typically reside in GPU memory. This secondary index can further reduce memory needs, especially when using LPs/SPDs that are likely to have many repeated values, but it is not necessary when using PPR scores since they tend to have distinct values.
Table 1. Complexity comparison of GRL models. Suppose using $O(|\mathcal{E}|)$-many queries, SGRLs use partial edges ($q\ll|\mathcal{E}|$) for training. $S$ and $K$ denote the average size of extracted subgraphs and sampled node sets, respectively. $L$ is the number of layers. $d$ and $k$ are respective dimensions of node and structural features. Assume $d$ is fixed for all layers. Both SUREL and SUREL+ use the walk-based sampler for $M$-many $m$-step walks. $c$ is the number of distinct $k$-dim structural features. $\delta_{n+1}$ is the size sum of all node sets.
Methods GNN (Kipf and Welling, 2017) SEAL (Zhang and Chen, 2018; Zhang et al., 2021) SUREL (Yin et al., 2022) SUREL+
Structure $O(|\mathcal{V}|+|\mathcal{E}|)$ $O(S|\mathcal{E}|)$ $O(mM|\mathcal{V}|)$ $O(\delta_{n+1})$
Feature $O(d|\mathcal{V}|)$ $O(kS|\mathcal{E}|)$ $O(\delta_{n+1}*k)$ $O(\delta_{n+1}+c*k)$
Time $O(|\mathcal{E}|Ld+|\mathcal{E}|Ld^{2})$ $O(qS^{L}d^{2})$ $O(qmMd^{2})$ $O(qKd^{2})$
Regarding the cost of SpG, indptr array is of size $|\mathcal{V}|+1$, and the size of both indices and SFptr arrays is $\delta_{n+1}$. The compressed encoding array $D_{\text{SF}}$, has a size of $c*k$, where $c$ is the number of distinct structural features and $k$ denotes feature dimension. The overall complexity of this data structure is $O(|\mathcal{V}|+\delta_{n+1}+c*k)$.
Comparison with Other Methods Table 1 summarizes the space and time complexity comparison of GRL methods. By adopting the walk-based sampler (sampling $M$-many $m$-step walks), $\delta_{n+1}$ amounts to around one-fifth space used by SUREL of $O(mM|\mathcal{V}|)$, while the metric-based sampler (sampling top-$K$ PPR scores) results in $\delta_{n+1}=K|\mathcal{V}|$ and $K<mM$ in general. Both values are substantially lower than $O(S|\mathcal{E}|)$ used by SEAL, where $S$ is the average size of sampled subgraphs. For hosting structural features, SUREL+ employs the secondary index SFptr $D$ and retains only distinct structural features in $D_{\text{SF}}$ to further reduce the memory footprint. In practice, $c$ typically remains independent of $|\mathcal{V}|$. Equipped with SpG, SUREL+ is capable of handling industry-level graph data.
3.3. Joining Node Sets via Sparse Operations
The goal of joining node sets is to connect queried nodes and construct query-induced subgraphs from pre-sampled node sets to make predictions for SGRL queries. For a given query $Q$, we merge relevant node sets $S_{u},\forall u\in Q$ into $\mathcal{S}_{Q}=\bigcup_{u\in Q}\mathcal{S}_{u}$ and join their node-level structural features $\mathcal{Z}_{u}$ to the query level. In essence, query-level structural features $\mathcal{Z}_{Q}$ serve as an anchor that indicates the position of node $x\in\mathcal{S}_{Q}$ within the query-induced subgraph $\mathcal{G}_{Q}$. Specifically, for a node $x$ in $\mathcal{S}_{Q}$, the query-level structural features $\mathcal{Z}_{Q,x}$ are obtained by joining its node-level structural features $\mathcal{Z}_{u,x}$ for all queried node $u$ in $Q$ as
(1) $\mathcal{Z}_{Q,x}=||_{u\in Q}\mathcal{Z}_{u,x}=[\dots\mathcal{Z}_{u,x}\dots]\in\mathbb{R}^{|Q|\times k},$
where ||||| | denotes concatenation. In cases where $\mathcal{Z}_{u,x}$ does not exist as node $x\notin S_{u}$ while $u\in Q$, it is set to all zeros. For instance, in Fig. 5, node $b$ is in $\mathcal{S}_{v}$ but not in $\mathcal{S}_{u}$, hence $\mathcal{Z}_{u,b}$ is set to zero. $\mathcal{Z}_{Q}$ is a collection of $\mathcal{Z}_{Q,x},\forall x\in\mathcal{S}_{Q}$. Together, $\mathcal{S}_{Q}$ and $\mathcal{Z}_{Q}$ function as the query-induced subgraph $\mathcal{G}_{Q}$, which is fed into the neural encoder later to obtain embeddings for prediction.
The JOIN operator in databases is used to merge tables and establish connections. Concatenation in Eq. (1) requires matching among different node sets with varying sizes and arbitrary node orders, where an outer JOIN is well-suited for this task. In this case, the JOIN operator returns associated values based on the node index (specified common field) from target sets, regardless of its existence. To obtain $\mathcal{Z}_{Q,x}$, node sets $\{\mathcal{S}_{u}\}_{u\in Q}$ are treated as tables: if the index of a node $x$ matches one of the node indices in $\mathcal{S}_{u}$ for all $u$ in $Q$, then the associated structural features $\mathcal{Z}_{u,x}$ are appended; otherwise, the field is filled with zeros. However, iterating over all $\mathcal{S}_{u}$’s to retrieve $\mathcal{Z}_{u,x}$ for each node $x\in\mathcal{S}_{Q}$ is highly inefficient, as its complexity can be $O(|Q|*|\mathcal{S}_{Q}|^{2})$ per query. This becomes even more challenging when performing these operations for massive queries with varying sizes of $\mathcal{S}_{u}$ and $\mathcal{S}_{Q}$.
To tackle this issue, we design an efficient arithmetic operator SpJoin to perform joins on sparse data objects of SpG in parallel. This operator reduces the time complexity to $O(|Q|*|\mathcal{S}_{Q}|)$ per query by taking advantage of node indices of $\mathcal{S}_{u}$ stored in SpG are unique and in sorted order. The following demonstrates the use of SpJoin for a query $Q=\{u,v\}$.
Sparse Join Operator The operator SpJoin performs an outer JOIN on the sampled node sets from seeds $u$ and $v$ of query $Q$ stored in SpG as $\texttt{SpG}[u,:]$ and $\texttt{SpG}[v,:]$ through
SpJoin $(\texttt{SpG}[u,:],\texttt{SpG}[v,:])=$
${\color[rgb]{0,0,1}\texttt{SpAdd}}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\texttt{mask},\texttt{SpG}[u,:]\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}{\color[rgb]{1,0,0}-1}~{}||~{}{\color[rgb]{0,0,1}\texttt{SpAdd}}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\texttt{mask},\texttt{SpG}[v,:]\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}{\color[rgb]{1,0,0}-1}$
where $\texttt{mask}={\color[rgb]{1,.5,0}\texttt{bool}}(\texttt{SpAdd}(\texttt{SpG}[u,:],\texttt{SpG}[v,:]))$.
As illustrated in Fig. 5, SpJoin consists of three steps:
  1. (1)
    It utilizes sparse arithmetic operations from SciPy (Virtanen et al., 2020): SpAdd performs an element-wise addition ($X\oplus Y$) of the non-zero elements in $X$ and $Y$; the resulting values are converted to binary via the bool operator and saved in the mask, which corresponds to node indices of the set $\mathcal{S}_{Q}$.
  2. (2)
    SpAdd are applied between mask and each $\texttt{SpG}[u,:],\,\forall u\in Q$ following by the reduction ‘-1’, which explicitly adds missing values (all zeros at default) to structural features $\mathcal{Z}_{u,x}$ for all $x$ if $x\not\in\mathcal{S}_{u}$ while $x\in\mathcal{S}_{v}$, for some $v\in Q\backslash\{u\}$.
  3. (3)
    When the secondary index SFptr $D$ is used, the results of SpJoin are pointers saved in SFptr, which can be used to gather the values of structural features $\mathcal{Z}_{Q}$ from the array $D_{\text{SF}}$.
Multithreading is employed to leverage the single program multiple data pattern in sparse operations of SpJoin. Since the processing time of each query linearly depends on the size of $\mathcal{S}_{Q}$, we further divide queries of each training batch into groups with nearly balanced sums of $|\mathcal{S}_{Q}|$’s, and assign one thread process a group to mitigate potential delays caused by uneven workloads.
Figure 5. An illustration of joining structural features from node-level to query-level (Eq. (1)) via the SpJoin operator on node sets stored in SpG. Note that SpG objects do not physically carry 00 as above shown in $\texttt{SpG}[u,:]$ and $\texttt{SpG}[v,:]$. Only non-zero elements in grey blocks are stored in SpG and performed arithmetic operations by SpJoin. The half-grey blocks correspond to added missing values.
Comparison with SUREL (Yin et al., 2022) SUREL adopts hash-based join operators to build query-level structural features, but its overall computation and memory cost is much higher compared to SUREL+. This is due to the presence of numerous repeated nodes in walks (over half as Fig. 3 (b) shown). The set-based input of SUREL+ substantially reduces the workload of transferring data from CPU to GPU and also requires fewer per-node operations on GPU to process transmitted $\mathcal{Z}_{Q}$’s. As Table 1 shows, SUREL+ reduces time complexity from $O(mM)$ to $O(K)$, where $K<mM$ is the average size of sampled node sets. These advantages ultimately enable SUREL+ to achieve superior efficiency and scalability.
3.4. Set Neural Encoders
After joining sampled node sets for each query $Q$, the resulting $(\mathcal{S}_{Q},\mathcal{Z}_{Q})$ acts as the query-induced subgraph $\mathcal{G}_{Q}$ and then is fed into neural encoders for prediction. The mini-batch training procedure of multiple queries is summarized in Algorithm 1. Next, we introduce neural encoders supported by SUREL+.
The adopted neural encoders are simple. For each $(\mathcal{S}_{Q},\mathcal{Z}_{Q})$,
(2) $h_{Q}=\texttt{AGGR}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\left(\color[rgb]{0,0,0}\{enc(\mathcal{Z}_{Q,x})|x\in\mathcal{S}_{Q}\}\definecolor{temp}{rgb}{0,0,0}\color[rgb]{0,0,0}\right)\color[rgb]{0,0,0}\in\mathbb{R}^{d}.$
Here, $enc(\cdot)$ encodes query-level structural features $\mathcal{Z}_{Q,x}$ using a multi-linear perception. If node attributes are present, they can be appended to the structural features as $\mathcal{Z}_{Q,u}||X_{u}$. AGGR is used to aggregate the encoded features, which can be any neural encoders applied on sets such as mean/sum/max pooling or set transformers. Currently, SUREL+ supports the implementations of AGGR in mean pooling, LSTM (Hamilton et al., 2017a), and attention (Veličković et al., 2018). Note that the LSTM applies random permutations to the elements in the set before encoding them as a sequence; while the attention first computes soft attention scores based on the output of $enc(\cdot)$ for each set element and then performs attention-score-weighted pooling. Sec. 4.4 empirically demonstrates that the choice of AGGR has non-trivial effects on prediction performance. Lastly, a fully connected layer takes the readout $h_{Q}$ as input to make the final prediction $\hat{y}_{Q}$. In our experiments, all SGRL tasks are formulated as binary classification, and thus Binary Cross Entropy is used as the loss function $\mathcal{L}$.
Input: Given a graph $\mathcal{G}(\mathcal{V},\mathcal{E},X)$, a group of queries $\{(Q,y_{Q})\}$ for training, batch size $B$, a set SAMPLER, a structure ENCODER, and a set AGGR
Output: A neural network for encoding subgraphs $enc(\cdot)$
1 Preprocessing: SAMPLER and ENCODER $\to(\mathcal{S}_{u},\mathcal{Z}_{u})$ for all $u\in\mathcal{V}$; convert and save $(\mathcal{S}_{u},\mathcal{Z}_{u})$’s as SpG objects.
2 for each mini-batch $\mathcal{Q}_{B}=\{...,Q,...\}$ do
3       Generate negative training queries (if not given) $\{\bar{Q}_{B}\}$ by random sampling and put them into $\mathcal{Q}_{B}$;
4       Call SpJoin operator to perform joining on SpG objects $\{(\mathcal{S}_{u},\mathcal{Z}_{u})|u\in Q\}$ for all queries $Q\in\mathcal{Q}_{B}$ in parallel;
5       Encode the joined results $(\mathcal{S}_{Q},\mathcal{Z}_{Q})$ as proxy of subgraphs via Eq. (2) with specified AGGR and get the prediction $\hat{y}_{Q}$ from readout $h_{Q}$ by multithreads;
6       Backward propagation based on the loss $\mathcal{L}(\hat{y}_{Q},y_{Q})$.
7 end for
Algorithm 1 The mini-batch training pipeline of SUREL+
4. Evaluation
In this section, we aim to evaluate the following questions:
  • Regarding space and time complexity, how much improvement can SUREL+ achieve by adopting sets instead of walks compared to the SOTA SGRL framework SUREL?
  • Can SUREL+ provide comparable prediction performance to all baselines using or not using SGRL methods?
  • How sensitive is SUREL+ to different choices of set samplers, structure encoders, and set neural encoders?
  • How do sparse storage SpG and parallelism in SpJoin operator benefit the overall performance of SUREL+?
4.1. Experiment Setup
Extensive experiments have been performed to evaluate SUREL+ using homogeneous, heterogeneous, and higher-order homogeneous graphs on three types of tasks: link prediction, relation type prediction, and higher-order pattern prediction. A homogeneous graph is a graph that does not include node/link types, while a heterogeneous graph includes various node/link types. Higher-order graphs are hypergraphs in our setting that contain hyperedges connecting two or more nodes.
Datasets Table 2 summarizes the statistics of datasets used to benchmark SGRL methods. Five datasets are selected from the Open Graph Benchmark (OGB, (Hu et al., 2020)) for link and relation type prediction, including social networks of citation - citation2 and collaboration - collab; biological network of protein interaction - ppa and blood vessels - vessel; and one heterogeneous academic network ogb-mag, which contains node types of paper (P), author (A) and their extracted relations. The vessel dataset is a large ($>$3M nodes), sparse, biological graph recently constructed from mouse brains (Paetzold et al., 2021) and has unique significance for examining GRL in scientific discovery. The structure of vessels illustrates the spatial organization of the brain’s microvasculature, which can be used for early detection of neurological disorders, e.g. Alzheimer’s and stroke. Two hypergraph datasets collected by (Benson et al., 2018) are used for higher-order pattern prediction: DBLP-coauthor is a temporal hypergraph, where each hyperedge denotes a time-stamped paper connecting all its authors. tags-math contains sets of tags applied to questions on the website math.stackexchange.com, represented as hyperedges. For higher-order pattern prediction tasks, the number of hyperedges is the main computation bottleneck, in which one may connect more than two nodes. Two industry-level graphs, criteo-click with 16.5M records of online banner ads clicking (Diemert et al., 2017) and twitter-2010 with 1.5B following relations of users (Kwak et al., 2010) are used to examine the model scalability for real-world applications.
Settings For link prediction, OGB’s standard data split is used to isolate validation and test links from the input graph. For prediction tasks of relation type and higher-order pattern, the same procedure to prepare graph data is adopted as in (Yin et al., 2022): the relations of paper-author (P-A, ”written by”) and paper-paper (P-P, ”cited by”) are selected; higher-order queries in hypergraph datasets are node triplets, where the goal is to predict whether it will foster in a hyperedge given two of them have observed pairwise connections; to learn the representation on hypergraphs, we project hyperedges into cliques and treat the projection results as ordinary graphs. All experiments are run 10 times independently, and we report the mean performance and standard deviation.
Table 2. Summary Statistics for Evaluation Datasets.
Dataset Type #Nodes #Edges Split(%)
criteo-click Homo./Bipartite
Campaign: 675
User: 6,142,256
16,468,027 97/1.5/1.5
twitter-2010 Homo./Social. 41,652,230 1,468,364,884 99.98/0.01/0.01
citation2 Homo./Social. 2,927,963 30,561,187 98/1/1
collab Homo./Social. 235,868 1,285,465 92/4/4
ppa Homo./Bio. 576,289 30,326,273 70/20/10
vessel Homo./Bio. 3,538,495 5,345,897 80/10/10
ogb-mag Hetero.
(P): 736,389
(A): 1,134,649
P-A: 7,145,660
P-P: 5,416,271
99/0.5/0.5
tags-math Higher. 1,629
projected: 91,685
hyperedges: 822,059
60/20/20
DBLP-
coauthor
Higher. 1,924,991
projected: 7,904,336
hyperedges: 3,700,067
60/20/20
Table 3. Prediction Performance for Links, Relation Types and Higher-Order Patterns: the best (bold) and the second best (underlined).
Models citation2 click twitter collab ppa vessel Models MAG(P-A) MAG(P-P) tags-math DBLP-coauthor
MRR (%) Hits@50 (%) Hits@100 (%) ROC-AUC MRR (%)
GCN 84.74±0.21 5.31±0.17 OOM 44.75±1.07 18.67±1.32 43.53±9.61 H*GCN 39.43±0.29 57.43±0.30 51.64±0.27 37.95±2.59
GraphSAINT 79.85±0.40 2.86±0.63 4.12±0.73 53.12±0.52 3.83±1.33 47.14±6.83 H*SAGE 25.35±1.49 60.54±1.60 54.68±2.03 22.91±0.94
GDGNN 86.96±0.28 13.30±0.45 49.86±0.39 54.74±0.48 45.92±2.14 75.84±0.08 R-GCN 37.10±1.05 56.82±4.71 - -
SEAL 87.67±0.32 OOM OOM 63.64±0.71 48.80±3.16 80.50±0.21 SUREL 45.33±2.94 82.47±0.26 71.86±2.15 97.66±2.89
SUREL 89.74±0.18 40.39±0.61 OOM 63.34±0.52 53.23±1.03 86.16±0.39 SUREL+ 58.81±0.42 80.45±0.13 77.73±0.16 99.83±0.02
SUREL+ 88.90±0.06 60.87±0.15 55.67±0.67 64.10±1.06 54.32±0.44 85.73±0.88 / / / / /
Table 4. Breakdown of Runtime, Memory Consumption for Different Models on Prediction of Link, Relation Type, and Higher-order Pattern. The column Train records the runtime per 10K queries.
Models Runtime (s) Memory (GB) Runtime (s) Memory (GB) Runtime (s) Memory (GB) Runtime (s) Memory (GB)
Prep. Train Inf. RAM SDRAM Prep. Train Inf. RAM SDRAM Prep. Train Inf. RAM SDRAM Prep. Train Inf. RAM SDRAM
Dataset criteo-click twitter-2010 citation2 ppa
GCN 3 0.085 8 3.1 62.74 - - - - OOM 17 21.74 105 9.3 36.84 2 0.026 1.2 4.6 11.35
GraphSAINT 1 0.012 20 13.1 8.79 111 0.009 920 253 76.60 151 1.79 107 9.6 9.78 10 0.003 1.5 4.9 23.06
GDGNN 215 1.43 2,928 16.2 23.77 1204 1.84 9,744 188 79.34 338 2.26 5,460 40.6 16.96 127 1.77 902 21.1 10.27
SEAL - - - OOM - - - - OOM - 46 3.52 24,626 35.4 5.71 46 10.57 3,988 9.5 12.13
SUREL 2 1.59 2,307 11.7 16.25 - - - OOM - 151 4.14 6,081 25.1 9.68 31 2.68 1,429 13.6 31.01
SUREL+ 22 0.23 502 10.4 11.93 327 0.26 3,779 210 49.44 130 0.35 1,389 16.7 4.75 69 0.72 201 9.8 19.02
Baselines We consider two classes of baselines. Canonical GNNs: GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017a), GraphSAINT (Zeng et al., 2020) and their variants with the prefix ‘H*’ that are directly applied for heterogeneous graphs with node types and for hypergraphs through clique expansion. R-GCN (Schlichtkrull et al., 2018) performs relational message passing on heterogeneous graphs; Scalable GNNs: and Cluster-GCN (Chiang et al., 2019); SGRL models: SEAL (Zhang and Chen, 2018; Zhang et al., 2021), GDGNN (Kong et al., 2022), and SUREL (Yin et al., 2022). SEAL adopts online subgraph sampling due to its intractable space needs for offline subgraph extraction. Fig. 3 (a) compares the time cost for subgraph sampling across different SGRL methods. We use all baselines’ official implementations with hyperparameter tuning to match their reported results.
Hyperparameters By default, SUREL+ uses the walk-based sampler, the structural encoder LP and the better set neural encoder tuned between mean pooling and attention. SUREL+ adopts a 2-layer MLP as $enc(\cdot)$ in Eq. (2) followed by a 2-layer MLP classifier to map set-aggregated representations for final predictions. Default training hyperparameters: learning rate lr=1e-3 with the early stopping of 5 epochs, dropout p=0.1, Adam (Kingma and Ba, 2015) as the optimizer. Analysis of parameters $M$ and $m$ to control the walk-based sampler and $K$ to control the metric-based sampler and selection of structure encoders and set neural encoders are studied in Sec. 4.4.
Evaluation Metrics The evaluation metrics include Hits@P, Mean Reciprocal Rank (MRR), and Area Under Curve (ROC-AUC). Hit@P counts the ratio of positive samples ranked at the top-P place against negative ones. MRR first computes the inverse of the rank of the first correct prediction and then takes the average of obtained reciprocal ranks for a sample of queries. For all datasets adopting MRR, each positive query is paired with 1000 randomly sampled negative test queries, except tags-math using 100 and crieo-click using 650. ROC-AUC follows the standard definition to measure the model’s performance in binary classification.
Environment We use a server with two Intel Xeon Gold 6248R CPUs, 512GB DRAM, and NVIDIA A100 (80GB) GPU. SUREL+ is built on PyTorch 1.12 and PyG 2.2. Set samplers are implemented in C, OpenMP, NumPy, Numba, and uhash, integrated into Python scripts; SpG is customized based on the CSR format of Scipy (Virtanen et al., 2020).
4.2. Prediction Accuracy Comparison
Table 3 shows the prediction performance of different methods. SGRL models significantly outperform canonical GNNs on these six link prediction benchmarks, especially on two challenging biological datasets ppa and vessel. Predicting links in biological datasets requires richer structural information that canonical GNNs have limited expressive power to capture. Within SGRL models, SUREL+ achieves comparable performance to SUREL and outperforms SEAL, which validates the effectiveness of the proposed set-based subgraph representation. For predictions of relation type and higher-order pattern, we observe additional performance gains (+2$\sim$13%) from SUREL+ compared to SUREL on three of the four datasets. A large performance gap exists between canonical GNNs and SGRL models, particularly in the higher-order case. This demonstrates the inherent limitations of canonical GNNs to make predictions of complex relations involving multiple nodes.
4.3. Efficiency and Scalability Analysis
Improved Efficiency in Training and Inference.
Table 4 compares model runtime and memory usage on the four largest benchmarks. SUREL+ offers a reasonable training time compared with canonical GNNs. It shows clear improvement in inference compared to the current SOTA framework SUREL (3-11$\times$ speedups across all datasets) and its predecessor SEAL ($\sim$20$\times$ speedups). SUREL+ achieves comparable and even lower RAM usage than canonical GNNs. Compared to other SGRL models, it can save up to half of RAM with lower usage of GPU DRAM, since set-based subgraphs eliminate node duplicates and their associated structural features. Tables 1 and 4 analytically and empirically show that the key factor of SUREL+ scaling up to billion-size graphs is its set-based representation with the sparse design, while GCN (full adjacency matrix), SEAL (complex subgraph extraction) and SUREL (dense walks with duplicate nodes) are all out of memory (OOM) on twitter-2010.
Profiling Different Strategies for Offline Processing
Fig. 5(a) reports the time cost of different samplers with multithreading on citation2. Fig. 5(b) shows memory consumption to store different types of sampled data (walks in SUREL (Yin et al., 2022) or sets in SUREL+) and their associated structural features (LPs, SPDs, PPR scores). Compared to SUREL sampler(Yin et al., 2022), the walk-based sampler in SUREL+ is more scalable and only adds one extra minute for encoding and converting data in SpG format (slash/dash marked in Fig. 5(a)), while achieving 6.94×6.94$\times$6.94 ×, 3.63×3.63$\times$3.63 × and 4.12×4.12$\times$4.12 × memory savings on three OGB datasets, respectively. Those memory savings are crucial for model scalability as they reduce the workload of data transfer from CPU to GPU and save GPU operations from encoding duplicated nodes, which dominates the time cost of the online stage and leads to the efficiency improvement of SUREL+ in Table 4. In addition, the metric-based sampler that adopts PPR scores has better scaling performance with more threads. When PPR scores or SPDs are adopted as structural features, SUREL+ further reduces the memory cost, though they often harm prediction performance slightly.
Note that in the above comparison of memory cost, compressed structural features are adopted both in SUREL (locally) and SUREL+ (globally), i.e., the secondary index based on SFptr $D$ and $D_{\text{SF}}$, which achieve compression of 493×493$\times$493 ×, 11318×11318$\times$11318 ×, 19527×19527$\times$19527 × on three datasets listed in Fig. 5(b), when one adopts LPs as structural features.
Scaling Analysis for SpJoin
Fig. 7 shows the speedups and throughput of the SpJoin operator for constructing query-level structural features via multithreading, where the join operation in SURE for nodes in walks (Yin et al., 2022) is used for comparison. SUREL employs a hash-based search for walk joins, which has unfavorable memory access patterns and suffers from imbalanced workloads due to inconsistent searching times across threads. SUREL+ gains more benefits from multithreading, thanks to sparse arithmetic operations and batch-wise load balancing used in SpJoin.
(a) Runtime
(b) Memory
Figure 6. Comparison of Runtime, Memory Consumption across Different Offline Processing Strategies (the walk-based sampler: $m=4,M=200$, the metric-based sampler: $K=150$). The areas highlighted break down the total consumption w.r.t. (a) sampling, structure encoding, sparse object construction; (b) structural features, node indices/pointers, and sampled walks (SUREL sampler only).
(a) Speedup
(b) Throughput
Figure 7. Scaling Performance Comparison of SpJoin in SUREL+ (with average set size $\bar{|S_{u}|}=351$) and Join Walks in SUREL (with walk size $m=4,M=200$) against Different Numbers of Threads.
4.4. Comparison between Different Set Samplers, Structural Features and Set Neural Encoders
SUREL+ is a modularized framework that supports different set samplers (walk- and metric-based), structural features (LP, SPD, PPR), and set neural encoders AGGR (mean pooling, LSTM, attention).
Table 5 shows the prediction performance and inference runtime by adopting different combinations of structure encoders and set neural encoders. Landing probabilities (LPs) as structural features perform the best on all three OGB datasets while being the slowest for inference. By recording the landing probabilities over different steps of walks, LPs provide structural information in finer granularity than scalar values of SPDs and PPR scores. The adopted link prediction task might favor more local information held by LPs and SPDs than global information carried by PPR scores. The authors conjecture that other tasks that rely on more global information may favor PPR scores. In comparison, no set neural encoder is always a winner. Attention seems to perform the best on average while slower than mean pooling. LSTM is the slowest. On the two social networks (citation2 and collab), mean pooling can provide comparable prediction results with much fewer parameters. However, prediction on the biological network (ppa) requires more expressive and complicated encoders, where LSTM and attention are favored as they can model more complex interactions between nodes in the joint set $\mathcal{S}_{Q}$.
Fig. 8 compares prediction results and inference time by using different hyperparameters $m,M$, and $K$ of set samplers, which heavily affects the coverage of sampled neighborhoods and computation overhead. The performance consistently increases if the walk-based sampler uses a larger $M$, but is not guaranteed for a larger $m$ (broader exploration). Better coverage with a larger $K$ is usually beneficial for the metric-based sampler over citation2 but not for collab, which is due to different characteristics of these two datasets and is also observed by (Yin et al., 2022). In general, small sampling parameters $m~{}(2\sim 4),M~{}(100\sim 400)$ and $K~{}(50\sim 200)$ can yield satisfactory performance with fast inference speed that achieves the trade-off between accuracy and scalability.
Table 5. Prediction Performance and Inference Runtime of SUREL+ with Different Combinations of Structure Features (LP, SPD, PPR) and Set Neural Encoders (Mean, LSTM, Attn.). The best and the second best are highlighted in bold and underlined accordingly.
Dataset PPR+Mean SPD+Mean LP+Mean LP+LSTM LP+Attention
citation2 78.59±0.38 87.99±1.07 88.55±0.15 88.46±0.34 88.90±0.06
834 1057s 1389s 3678s 2171s
collab 47.15±0.21 62.11±0.13 64.10±1.06 61.31±1.37 62.85±1.19
1.4s 1.7s 2.0s 3.5s 2.3s
ppa 13.28±1.20 41.06±1.70 46.41±1.65 54.45±1.35 54.32±0.44
63s 126s 165s 322s 201s
Figure 8. Hyperparameter Analysis of Set Samplers (Prediction Performance v.s. Time Cost). Walk-based: the number $M$ and the step $m$ of walks, LPs as structural features; Metric-based: the set size $K$, PPR scores as structural features.
5. Conclusion
In this work, we propose a novel framework SUREL+ for scalable subgraph-based graph representation learning. SUREL+ avoids costly query-specific subgraph extraction by decoupling it into sampled node sets with structural features whose join can function as query-induced subgraphs for prediction. SUREL+ benefits from the reusability and compactness of pre-sampled node sets across different queries. Compared to the SOTA SGRL framework SUREL, the set-based representation of SUREL+ substantially reduces space and time complexity by avoiding heavy node duplicates in sampled walks. To handle irregularly sized node sets, SUREL+ designs a customized sparse storage SpG and a sparse join operator SpJoin, providing memory-efficient storage with fast access. In addition, SUREL+ adopts a modular design, enabling users to choose different set samplers, structure encoders, and set neural encoders flexibly based on the nature of their SGRL tasks. Extensive experiments on three types of prediction tasks over nine real-world graph benchmarks show that SUREL+ significantly improves scalability, memory efficiency, and prediction accuracy compared to current SGRL methods and canonical GNNs.
Acknowledgements.
The authors would like to thank Rongzhe Wei and Yanbang Wang for their helpful discussions and valuable feedback. Haoteng Yin and Pan Li are supported by the 2021 JPMorgan Faculty Award, NSF awards OAC-2117997, IIS-2239565.
Appendix A Notations
Frequently used symbols are summarized in Table 6.
Table 6. Summary of Frequently Used Notations.
Symbol Meaning
Q
a query (set of nodes), i.e. $Q=\{u,v,w\}$
$\mathcal{Q}$
a set of queries, i.e. $Q\in\mathcal{Q}$
$\mathcal{G}_{u}$
a subgraph induced by node $u$
$\mathcal{G}_{Q}$
a subgraph induced by query $Q$
$\mathcal{S}_{u}$
a set of unique nodes sampled from the neighborhood of the seed node $u$
$\mathcal{Z}_{u,x}$
structural features of node $x$ regarding the seed node $u$ (all zeros if $x\notin\mathcal{S}_{u}$)
$\mathcal{Z}_{u}$
collection of structural features for all nodes in $\mathcal{S}_{u}$ as $\mathcal{Z}_{u}=\{\mathcal{Z}_{u,x}|x\in\mathcal{S}_{u}\}$
the concatenation that joins node-level structural features, i.e. join $\mathcal{Z}_{\cdot,x}$ for a query $Q=\{u,v,w\}$ as $[\mathcal{Z}_{u,x},\mathcal{Z}_{v,x},\mathcal{Z}_{w,x}]$.
$\mathcal{Z}_{Q,x}$
query-level structural features for node $x$ regarding the query $Q$, $\mathcal{Z}_{Q,x}=||_{u\in Q}\mathcal{Z}_{u,x}$
Table 7. Summary Statistics and Experimental Setup for Evaluation Datasets.
Dataset Type #Nodes #Edges Avg. Node Deg. Density Split Ratio Split Type Metric
criteo-click Homo./Bipartite
Campaign(C): 675
User(U): 6,142,256
16,468,027 2.68 N/A 97/1.5/1.5 Time MRR
twitter-2010 Homo./Social. 41,652,230 1,468,364,884 35.25 0.00017% 99.98/0.01/0.01 Random MRR
citation2 Homo./Social. 2,927,963 30,561,187 20.7 0.00036% 98/1/1 Time MRR
collab Homo./Social. 235,868 1,285,465 8.2 0.0046% 92/4/4 Time Hits@50
ppa Homo./Bio. 576,289 30,326,273 73.7 0.018% 70/20/10 Throughput Hits@100
vessel Homo./Bio. 3,538,495 5,345,897 3.02 0.000085% 80/10/10 Random AUC-ROC
ogb-mag Hetero.
Paper(P): 736,389
Author(A): 1,134,649
P-A: 7,145,660
P-P: 5,416,271
21.7 N/A 99/0.5/0.5 Time MRR
tags-math Higher. 1,629
91,685 (projected)
822,059 (hyperedges)
N/A N/A 60/20/20 Time MRR
DBLP-coauthor Higher. 1,924,991
7,904,336 (projected)
3,700,067 (hyperedges)
N/A N/A 60/20/20 Time MRR
Table 8. [Extended] Breakdown of Runtime, Memory Consumption for Different Models on Prediction of Link, Relation Type, and Higher-order Pattern. The column Train records the runtime per 10K queries.
Models Runtime (s) Memory (GB) Runtime (s) Memory (GB) Runtime (s) Memory (GB) Runtime (s) Memory (GB)
Prep. Train Inf. RAM SDRAM Prep. Train Inf. RAM SDRAM Prep. Train Inf. RAM SDRAM Prep. Train Inf. RAM SDRAM
Dataset citation2 ppa collab vessel
GCN 17 21.74 105 9.3 36.84 2 0.026 1.2 4.6 11.35 2 0.005 0.05 2.5 5.50 5 0.076 0.3 2.8 36.98
GraphSAINT 151 1.79 107 9.6 9.78 10 0.003 1.5 4.9 23.06 1 0.004 0.08 2.5 8.11 5 0.008 15 6.9 10.21
GDGNN 338 2.26 5,460 40.6 16.96 127 1.77 902 21.1 10.27 14 0.74 15 4.3 1.08 25 0.85 84 7.2 8.03
SEAL 46 3.52 24,626 35.4 5.71 46 10.57 3,988 9.5 12.13 5 4.05 37 4.0 6.20 6 10.69 998 6.2 2.46
SUREL 151 4.14 6,081 25.1 9.68 31 2.68 1,429 13.6 31.01 1 2.13 17 3.4 9.86 5 1.57 32 5.8 5.18
SUREL+ 130 0.35 1,389 16.7 4.75 69 0.72 201 9.8 19.02 7 0.27 2 2.8 3.37 3 0.31 3 3.3 1.25
Dataset MAG(P-A) MAG(P-P) tags-math DBLP-coauthor
H*GCN 3 0.03 9 5.0 21.56 4 0.03 13 5.5 21.66 2 0.004 1.3 2.4 3.10 - 0.58 95 8.0 25.80
H*SAGE 3 0.03 10 5.0 20.29 4 0.03 13 5.5 20.28 1 0.003 1.3 2.4 3.10 - 0.32 77 7.5 24.70
R-GCN 1 0.52 5 5.3 26.34 1 0.52 4 5.1 31.41 - - - - - - - - - -
SUREL 10 3.20 1,998 7.3 7.18 15 0.99 1924 8.1 16.66 - 2.13 341 3.0 5.95 11 1.29 949 9.7 7.79
SUREL+ 58 0.33 101 7.2 2.95 77 0.13 168 8.1 13.49 1 0.67 116 2.4 5.70 8 0.24 315 3.8 3.16
Appendix B More Details
B.1. Other Related Works
Scalable GNN Design. GNNs are the most widely used tools for graph representation learning nowadays, although they face certain challenges when directly applied to subgraph-based methods. To address the scalability of GNNs, current studies focus on improving graph subsampling and mini-batch training techniques (Chiang et al., 2019; Zeng et al., 2020). However, graph subsampling used in GNNs fundamentally differs from subgraph extractions in SGRL. The goal of subsampling is to handle GPU memory overflow during full-batch training of GNN models. For SGRL, subgraphs sampled around a query serve as features for making predictions. Consequently, the scaling techniques developed for GNNs cannot be directly applied to SGRL. Another direction is to deploy distributed GNN systems for industry-level graphs. Unfortunately, these specialized techniques, including pipelining (Wan et al., 2022b), partitioned parallelism (Wan et al., 2022a), and update with staleness (Peng et al., 2022) do not address the main bottleneck of subgraph extraction for SGRL methods.
B.2. Model Design
Benefits of Subgraph-based Graph Representation Learning First, subgraph-based representation is versatile for different types of tasks, especially when queries of certain tasks involving multiple nodes and relations, e.g. existence of a link, property of a motif, development of higher-order patterns; while canonical GNNs are limited to handle such polyadic dynamics via node-wise representations (Srinivasan and Ribeiro, 2020; Wang and Zhang, 2021) (also refers to the example in Fig. 2). Second, subgraph-based models are more expressive by pairing with structural features to obtain most expressive structural node representations (Srinivasan and Ribeiro, 2020; Li et al., 2020; Wang and Zhang, 2021; Bouritsas et al., 2022). However, canonical GNNs cannot capture intra-distance information and joint relations over multiple nodes, which are critical to distinguishing nodes in structural symmetry and making predictions over them. Lastly, subgraph-based methods decouple the model depth from the receptive field since extracted subgraphs are localized to certain hops: when adding more layers for non-linearity, it does not contaminate embedding with irrelevant nodes or get over-smooth as canonical GNNs do. This results in a more robust representation and is particularly beneficial for modeling relations beyond singleton.
B.3. Datasets
The full statistics of benchmark datasets are summarized in Table 7. OGB datasets [^1] https://ogb.stanford.edu/docs/dataset_overview/ are selected to benchmark our proposed framework and other baselines. It contains large-scale graphs (millions of nodes/edges) for real-world applications (i.e., academic and biological networks) and provides standard, open-sourced evaluation metrics and tools. Note that, vessel is a newly added benchmark of a biological graph, with $>3M$ nodes and sparse vessel structures extracted from the whole mouse brain (Paetzold et al., 2021), where nodes represent bifurcation points, and edges represent the blood vessels. Each node is associated with features of its physical location in the coordinate space $(x,y,z)$. The introduction of vessel provides a unique opportunity to examine graph representation learning (GRL) approaches in neuroscience, especially in scaling subgraph-based GRL methods to handle sparse and spatial graphs with millions of nodes and edges for scientific discovery.
criteo-click contains a sample of 30 days of Criteo live traffic data, each corresponding to one impression (a banner) displayed to a user and whether it is clicked (Diemert et al., 2017). Each record has 9 contextual features that are aggregated into a 270-dimensional edge feature. There are 675 unique campaign banners and 6.1M users, consisting of a bipartite graph of 16.5M edges: 97% is used for training, and the rest is evenly split for validation and testing based on temporal orders. The task is to predict which campaign the user is most likely to click among 651 candidates. twitter-2010 is an industry-level social network with 1.5B following relations of users (Kwak et al., 2010). An edge $(i,j)$ of this network indicates that user $i$ is followed by user $j$. 1% of Twitter users who follow 10 to 1000 accounts are randomly sampled for evaluation. The task is to recommend which account they will most likely follow among 1001 candidates. The OGB formatted files of these two datasets are accessible via Box at https://purdue.box.com/v/SGRL-LSC-dataset.
B.4. Baselines
For link prediction and relation type prediction, baseline models are selected based on their scalability and prediction performance from the current OGB leaderboard  [^2] https://ogb.stanford.edu/docs/leader_linkprop/. All models listed on the leaderboard are publicly accessible. We adopt their reported numbers on the leaderboard with verification. For the rest of the baselines, we benchmark these models using their official implementations with tuned hyperparameters as listed below.
  • Canonical GCNN: a graph auto-encoder model that uses graph convolution layers to learn node representations, including GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017a), and their more scalable variants by employing graph subsampling, such as Cluster-GCN (Chiang et al., 2019), GraphSAINT (Zeng et al., 2020).
  • R-GCN [^3] https://github.com/pyg-team/pytorch_geometric/blob/master/examples (Schlichtkrull et al., 2018): a relational GCN that models heterogeneous graphs with different types of node/link.
  • SEAL [^4] https://github.com/facebookresearch/SEAL_OGB (Zhang and Chen, 2018): apply GCN on exact query-induced subgraphs with double radius node labeling to obtain subgraph-level readout for link prediction. SEAL shows great empirical performance on multiple graph machine learning benchmarks and promotes the deployment of subgraph-based models for scientific discovery. The implementation we tested is specialized for OGB datasets provided in (Zhang et al., 2021).
  • GDGNN [^5] https://github.com/woodcutter1998/gdgnn (Kong et al., 2022): a model aggregates node representations generated by GNNs along geodesic paths between nodes in a query to speedup subgraph-based methods.
  • SUREL [^6] https://github.com/Graph-COM/SUREL(Yin et al., 2022): a walk-based computation framework to accelerate subgraph-based methods, where subgraphs are decomposed to pre-sampled walks and then are joined online to substitute the query-induced subgraph for prediction. SUREL achieves state-of-the-art scalability and prediction accuracy on SGRL tasks by adopting walk-based representation.
All canonical GNN baselines [^7] https://github.com/snap-stanford/ogb/tree/master/examples/linkproppred come with three GCNConv/SAGEConv layers of 256 hidden dimensions, and a tuned dropout ratio in $\{0,0.5\}$ for full-batch training. Canonical GNN models aggregate all node embeddings involved in a query as the representations of links/hyperedges, which are later fed into an MLP classifier for final prediction. In addition, all GNN models need to use full training data (edges/triplets) to generate robust node representations. The hypergraph datasets do not come with raw node features, thus GNN baselines here use randomly initialized features as input for training along with other model parameters. R-GCN uses RGCNConv layers that support message passing with multiple relation types between different types of nodes, where the edge types (relations) are used as input beside node features.
Subgraph-based models only use partial edges/triplets for training. For SEAL, 1-hop enclosing subgraphs are extracted online during the training and inference. Then, it applies three GCN layers of 32 hidden dimensions plus a sort pooling and several 1D convolution layers to generate a readout of the target subgraph for prediction. SUREL consists of a 2-layer MLP for relative position encoding (RPE) and a 2-layer RNN to encode joined walks attached with RPEs to obtain the readout of query-level subgraphs. The hidden dimension of both networks is set to 64. The obtained embeddings of joined walks are aggregated and fed into a 2-layer MLP classifier to make predictions. GDGNN employs GINLayer as its backbone to obtain node embeddings. The horizontal geodesic representation is used for predictions, which finds the shortest path between two nodes in a query and aggregates node representations generated by GNNs along the found geodesic path. The max search distance for geodesic is the same as the number of GNN layers. For collab, ppa, citation2 and vessel, the threshold of distance is set to 4, 4, 3, and 2, respectively. The hidden dimension of all fully connected layers is set to 32.
Table 9. Hyperparameters Used for Benchmark SUREL+.
Dataset #steps $m$walks $M$
#negative samples $k$
Structural Feature
Set Neural Encoder
criteo-click 4 200
10
LP
Mean
twitter-2010 4 100
25
LP
Mean
citation2 4 100
10
LP
Mean
collab 3 200
10
LP
Mean
ppa 4 200
20
LP
Attn.
vessel 2 50
5
LP
Mean
MAG (P-A) 3 200
10
LP
Mean
MAG (P-P) 4 100
10
LP
Mean
tags-math 4 200
10
LP
Mean
DBLP-coauthor 3 100
10
LP
Mean
Appendix C Architecture and Hyperparameter
SUREL+ uses a 2-layer MLP with ReLU activation for encoding structural features and supports three set neural encoders, including mean pooling, LSTM, and attention. LSTM interprets elements to be aggregated in a set as a sequence (Hamilton et al., 2017a); attention first calculates soft attention scores for elements in a set and then performs attention-score-weighted average pooling. The hidden dimension of all parameterized layers is set to 96. Lastly, hidden representations of joined node sets are fed into a 2-layer MLP classifier for final predictions.
For the implementation of set samplers, the walk-based sampler builds on the sampling function from SubGAcc [^8] https://github.com/VeritasYin/subg_acc library developed by the authors, which also provides the support for efficient structural feature compression and index remapping. The metric-based sampler is adopted from fast PPR approximation in (Bojchevski et al., 2020).
We follow the inductive setting for link and relation prediction: only partial samples will be used for training. Over the training graph, we randomly select 5% links as positive training queries, each paired with $k$-many negative samples ($k=10$ by default). We mask these links and use the remaining 95% links to compute each node’s structural features in the split training set via structure encoder. For vessel, as the input graph is very sparse, we first sort the nodes in training set by their degree and then randomly pick 5% nodes to obtain edges of their 2-hop induced subgraphs for training and the rest reserved for structural feature construction. For higher-order pattern prediction, we use the given graph before timestamp $t$ to sample node sets and encode their structural features. The model parameters are optimized by triplets provided in the training set. No node features are used in SUREL+, except for vessel where normalized physical locations of each node are attached after its structural features and contextual features of click are attached.
The extended version of Table 4 is provided in Table 8 for the comparison on the rest of the datasets. The results reported in Table 3 and the profiling of SUREL+ in Tables 4, 8 are obtained through the combination of hyperparameters listed in Table 9. The dropout rate on vessel is set to p=0.2. The metric-based sampler is adopted to obtain the results of using PPR and SPD as structural features in Table 5. Its sampling size $K$ is set to $50$, $50$ and $150$ for collab, ppa, citation2, respectively. The walk-based sampler is used for the results of LP as structural features, whose sampling parameters are listed in Table 9. The rest of the hyperparameters remain the same as reported in Sec. 4.1. The SUREL+ framework including SubGAcc library is open-source and free for academic use under the BSD-2-Clause license.
b9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">